OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2)

OpenStudy (eyust707):

\[(y - y_{1}) = m (x-x_1)\] m = 3 point = (2, 5)

OpenStudy (anonymous):

y – 5 = 3(x – 2)

OpenStudy (eyust707):



OpenStudy (eyust707):